# ActiView: Evaluating Active Perception Ability for Multimodal Large Language Models

### v1

@@ Abstract
Active perception, a crucial human capability, involves setting a goal based on the current understanding of the environment and performing actions to achieve that goal. Despite significant efforts in evaluating Multimodal Large Language Models (MLLMs), active perception has been largely overlooked. To address this gap, we propose a novel benchmark named ActiView [^1] Codes and data will be available at https://github.com/THUNLP-MT/ActiView to evaluate active perception in MLLMs. Since comprehensively assessing active perception is challenging, we focus on a specialized form of Visual Question Answering (VQA) that eases the evaluation yet challenging for existing MLLMs. Given an image, we restrict the perceptual field of a model, requiring it to actively zoom or shift its perceptual field based on reasoning to answer the question successfully. We conduct extensive evaluation over 27 models, including proprietary and open-source models, and observe that the ability to read and comprehend multiple images simultaneously plays a significant role in enabling active perception. Results reveal a significant gap in the active perception capability of MLLMs, indicating that this area deserves more attention. We hope that our benchmark could help develop methods for MLLMs to understand multimodal inputs in more natural and holistic ways.
11footnotetext: These authors contribute equally.
@@ 1 Introduction
The advent of Multimodal Large Language Models (MLLMs) has marked a significant milestone in the realm of artificial intelligence, demonstrating capabilities that are increasingly approaching human-like performance (OpenAI, 2023; Liu et al., 2023c; Ye et al., 2024b). This advancement, while promising, also presents new challenges and opportunities for evaluating these models. As a result, the landscape of MLLM evaluation is rapidly evolving, with numerous benchmarks being developed to either comprehensively evaluate models (Fu et al., 2023; Liu et al., 2023d; Fu et al., 2024a) or to analyze specific aspects of their capabilities (Liu et al., 2023a; Lu et al., 2023; Luo et al., 2024; Xiao et al., 2024; Li et al., 2024b; Nie et al., 2024; Qian et al., 2024).
Despite the extensive efforts devoted to MLLM evaluation, active perception (Bajcsy, 1988; Bajcsy et al., 2018) remains underexplored. Active perception involves understanding the reasons for sensing, choosing what to perceive, and determining the methods, timing, and locations for achieving that perception (Bajcsy et al., 2018). This is important because in the real world, the desired information often does not appear directly in the center of one’s field of vision. Instead, it requires individuals to move their field of view, locate details, and filter out distracting information. For example, in Figure 1(a), suppose we are looking for information in a giant painting. We need to first shift our view to locate the specific area and then possibly zoom in to gather detailed information. Intuitively, active perception not only enables a person or model to accomplish more complex tasks, but it also has the potential to serve as a good indicator of the level of intelligence of a model. This makes it a critical capability that warrants thorough evaluation.
However, existing multimodal evaluation benchmarks are not well-suited for measuring active perception capabilities. We summarize several widely used or recently proposed multimodal evaluation benchmarks in Table 1. Most of these benchmarks assess models in static perceptual field settings, evaluating how well a model can process information presented directly to it without requiring active exploration or dynamic adjustments to its field of view. BLINK (Fu et al., 2024b), V* (Wu & Xie, 2023), and CNT (Roberts et al., 2023) are the only exceptions, as they utilize dynamic perceptual fields. However, they only consider either shifting or zooming of the field of view in specific scenarios, which are insufficient for measuring active perception capabilities. Therefore, there is a clear need for new evaluation frameworks that can adequately assess active perception abilities across diverse and dynamic environments.
To fill this gap, we introduce a novel benchmark specifically designed for evaluating Active perception through View changes (ActiView). Given the difficulty of comprehensively evaluating active perception capabilities across all possible scenarios, ActiView concentrates on a series of tasks that are currently feasible to evaluate yet still present significant challenges to current models. The evaluation instances in ActiView follow the Visual Question Answering (VQA) (Antol et al., 2015) format but include the following additional features: 1) Each question requires an understanding of detailed visual clues in the image to answer accurately. 2) We impose view constraints on images, allowing the model to perceive only a partial field of view of the full image at a time. This setup explicitly requires models to perform view shifting and zooming to gather more information and eliminate potential distractions, thereby simulating the active perception process. We collect images from existing datasets as well as newly captured images, carefully curate questions, answers, and visual clues manually to ensure the quality and diversity of the benchmark.
@@
Figure 1: (a) Active perception allows humans or models to perform more complex tasks by actively seeking and processing relevant information. In this paper, we evaluate two key active perception abilities for MLLMs: 1) shifting, as real-world scenarios often present limited views and require shifts to obtain new perspectives, and 2) zooming, which helps enhance perception by zooming out for broader views and zooming in for details. (b) Model results on our benchmark compared to human performance, showing that our benchmark is challenging for models.
We conducted evaluations on three proprietary models and 24 widely used and advanced open-source models. The results reveal that these models generally lag behind in active perception. For instance, the strong proprietary model, GPT-4o, only achieved an average score of 66.40% with our designed evaluation pipelines, and a score of 67.38% on the general VQA format, which is significantly lower than the human performance score of 84.67%. Moreover, the average performance gap between proprietary models and open-source models in active perception is considerably smaller within our designed evaluation pipelines compared to gaps observed in other tasks from previous research. Experimental results suggest that models tend to perform better when the full image is provided, while fall shot to form a holistic perspective of the complete image when they only have access to all the constrained perceptual fields. Therefore, we believe that active perception warrants more research efforts, and our benchmark is valuable for further study.
Table 1: Comparison with other benchmarks for MLLMs. “Per. Fields”: Perceptual Fields. 1.9k*: Videos. Manual*: A mixture of manual annotation and data from existing benchmarks. Our benchmark concentrates on evaluating active perception abilities via the change of visual perceptual fields, including shifting to different fields for compensating missing information, and zooming for fine-grained details in the current fields.
@@
MME (Fu et al., 2023) Visual comprehension 1.1k 1.3k Manual
@@
Video-MME (Fu et al., 2024a) Visual comprehension 0.9k 2.7k Manual
ViP-Bench (Biernacki et al., 2021) Understanding of visual prompt 303 303 Manual
V* (Wu & Xie, 2023)
Visual search for
fine-grained details
191 191 Manual
@@
BenchLMM (Cai et al., 2023) Cross-style visual comprehension 1.7k 1.7k Manual + Auto
CODIS (Luo et al., 2024)
Context-dependent
visual comprehension
377 754 Manual
@@
2 Related Work
@@ 2.1 MLLM Benchmarks
Extensive efforts have been devoted to developing MLLM evaluation benchmarks (Table 1). These benchmarks cover a wide range of capabilities, including but not limited to visual comprehension (Fu et al., 2023; Liu et al., 2023d; Fu et al., 2024a), visual perception (Fu et al., 2024b), hallucination (Liu et al., 2023a), and mathematical and logical reasoning (Lu et al., 2023; Xiao et al., 2024). However, most of these benchmarks utilize a static view of the input image, which is not suitable for evaluating active perception. While BLINK (Fu et al., 2024b) involves view shifting and both V* (Wu & Xie, 2023) and CNT (Roberts et al., 2023) require view zooming, active perception is not a prerequisite for solving their evaluation questions, making them insufficient for active perception evaluation. In contrast, our benchmark considers both view shifting and zooming. Questions in our benchmark are specifically designed to necessitate active perception for answering, making it a more comprehensive benchmark for active perception evaluation. Active perception is also used to study the control of model states involving dynamic interactions between models and environment (Oh et al., 2016; Ammirato et al., 2017). However, in this paper, we focus on how models actively perceive and understands visual information, which are orthogonal to these works.
@@ 2.2 Active Perception in MLLMs
Although MLLMs have attracted extensive interest, less effort has been dedicated to improving the active perception capability of MLLMs. One line of research focuses on improving the ability of processing high-resolution images by using higher-resolution ViTs (Ye et al., 2024b), slicing high-resolution images and then concatenate them (Liu et al., 2024), or directly using LLMs to process raw patches of any resolution (Li et al., 2023a). The other line emphasizes visual search for fine-grained details. For example, SEAL (Wu & Xie, 2023) fine-tunes a framework of two MLLMs to follow the visual search mechanism for precise visual grounding, and V-IRL (Yang et al., 2024) proposes an active detection strategy to improve the comprehension of real-world geospatial information. Despite these efforts, our evaluation results reveal that existing MLLMs still generally lack active perception capabilities. Our benchmark will shed light on evaluating and enhancing active perception in MLLMs.
@@ 3 ActiView
Our benchmark exams active perception abilities of models via different perceptual fields, where Actively zooming and shifting of Views (ActiView) are required. We summarise zooming and shifting as two fundamental factors of active perception as illustrated in Figure 2, where we can evaluate active perception abilities of models through the two factors separately or integratedly. ActiView imitates the procedure of active perception by providing models with an initial view, which is a cropped field of the original image, as depicted in Figure 2, and requires models to search for missing but important information with view zooming and shifting, and to minimize distractions caused by redundant information in the view. We manually annotated 325 questions, where each question corresponds to 5 different evaluation instances that assess active perception abilities under different settings and levels of difficulty. In total, there are 1,625 evaluation instances in our benchmark. This section elaborately describes ActiView benchmark, including the design of categories, questions and options (§3.1) and the data collection procedure and statistics of the benchmark (§3.2).
Figure 2: Examples of ActiView. Instances in ActiView present the following features: i) requiring focusing on multiple fine-grained regions; ii) requiring distinguishing distracting information from the entire image; iii) requiring moving of perceptual fields to obtain sufficient visual information to answer questions. During evaluation, models will be given an initial view cropped from the original image as shown above. The evaluation procedures given the initial views are introduced in §4.
@@ 3.1 Benchmark Overview
When perceiving an image, humans intuitively focus on three principle aspects: the environment depicted in the image, the primary objects, and the event that these objects are engaged in. Correspondingly, we summarise the questions in our benchmark into three main categorises, including environment-centric (Type I), object-centric (Type II), and event-centric (Type III) categorises. These main categorises are further divided into eight sub-classes according to the specific type of visual information and visual features used for answering the questions. Figure 2 displays eight examples for each of the sub-classes, and demonstrates features of our benchmark. For the environment-centric category, three sub-classes are developed:
  • Geo-Localization, denoted as Geo-Loc. This sub-class focuses on geographical features that are unique to a country or a city, and requires models to identify geographical locations depicted in target images. Typical questions for this sub-class are “Where is this place located?”,“In which country is the photo taken?”, and etc. Images in this sub-class usually contain unique landmarks such as the Eiffel Tower in Paris and the Atomium in Brussels.
  • Orientation, denoted as Orient. This sub-class challenges models to exploit natural orientation information for answering the questions, such as the position of shadows, the position of the sun, and the directional information on street signs. Questions of this type include “Is this a sunset or a sunrise?”, “Where is the sunlight coming from?” and etc.
  • Daily-location, denoted as Daily-Loc. To distinguish from the Geo-Loc sub-class, this sub-class concentrates on locations in everyday life that could appear in most of the cities and are not unique to a certain city or country. Images in this sub-class usually depict scenes of museums, restaurants, shops, etc. The corresponding questions include “Where is this picture most likely taken?”, “Is there a music school nearby?”, and etc.
  • Object-attribute, denoted as Obj-Attr. These questions ask for the attributes of objects, with the appearance of distracting information that potentially lead to incorrect answers. As the example of Obj-Attr shown in Figure 2, the highest price, 69 per kilogram, corresponds to papaya rather than watermelon.
  • Object-relation, denoted as Obj-Rel, which concentrates on the spatial relationships among multiple objects. Questions in this sub-class do not directly ask for the spatial relationships. They require models to reason for the correct answer via spatial information. Figure 2 displays an example of Obj-Rel in which models should be aware of the relative positions of the feet of the person to the water.
  • Counting, denoted as Count-Dis. Although this sub-class focuses on the number of objects, different from the counting tasks in other benchmarks (Fu et al., 2023; Yu et al., 2023), there are similar but distracting information about the targets in the images. These distracting objects easily confuse models and challenge the abilities to understand and strictly follow instructions. As the Count-Dis example shown in Figure 2, the jerky on the table should not be considered when answering the question of “How many pieces of jerky are hanging on the wall?”.
  • Event-single, denoted as Event-S. There is only one item or person involved in the target event. For example, the image for Event-S in Figure 2 shows one person driving without other people presenting in the image.
  • Event-multi, denoted as Event-M. Different from Event-S, events of this type happen among multiple items or people. As the example of Event-M shown in Figure 2, the “woman in blue” is engaged in a photo shooting activity in which she is posing and another person is taking photo for her. This type requires models to distinguish the event (or events) each entities presenting in the image are engaged in.
To prevent MLLMs from directly answering the questions by merely going through the options, we formulate most of candidate options from the images themselves. For instance, as the example of Geo-Loc shown in Figure 2, there are corresponding information for each of the options in the image, where flags representing UK, France, Spain and Hungary appear. Additionally, for options comprised by numbers, we display them in random order to avoid biased predictions.
3.2 Data Collection and Data Statistics
Our data are manually curated, including image collection and question annotation processes. To evaluate active perception abilities requiring zooming of fine-grained details and shifting views to obtain missing information, we carefully select images containing multiple fine-grained objects and depicting complex environment and events. To assure the clearness of visual details and the quality of initial views, we collect high-resolution images as described in Appendix A, together with annotation guidelines to ensure the fairness and reliability of question-option-answer triplets. Additionally, annotators are instructed to identify and highlight the corresponding visual clues in the image that support their answer. In Figure 2, the human-annotated visual clues are presented as sub-images placed below the original images in “Visual Information” columns. Furthermore, these visual clues will serve as criteria for automatic sorting of different difficulties. We also discuss approaches for automatic data generation in Appendix J. As a brief conclusion here, powerful models, such as GPT-4V, fail to satisfy our annotation criteria, because they suffer from hallucination when comprehending images, and fail to distinguish visual facts from the image and external world knowledge that does not present in the image.
Figure 3: The distribution of categories and sub-classes of our benchmark.
The distribution of categories are shown in Figure 3. We manually collected 314 images and annotated 325 questions. For each of the questions, the number of candidate options ranges from two to seven, with an average of 3.24 options per question. There are 2.64 sub-views containing human-annotated clues for each of the questions on average. This indicates that a single view is not sufficient for obtaining the correct answer, and that the ability to read and comprehend multiple images is indispensable to our benchmark.

### v2

@@ Abstract
Active perception, a crucial human capability, involves setting a goal based on the current understanding of the environment and performing actions to achieve that goal. Despite significant efforts in evaluating Multimodal Large Language Models (MLLMs), active perception has been largely overlooked. To address this gap, we propose a novel benchmark named ActiView to evaluate active perception in MLLMs. We focus on a specialized form of Visual Question Answering (VQA) that eases and quantifies the evaluation yet challenging for existing MLLMs. Meanwhile, intermediate reasoning behaviors of models are also discussed. Given an image, we restrict the perceptual field of a model, requiring it to actively zoom or shift its perceptual field based on reasoning to answer the question successfully. We conduct extensive evaluation over 30 models, including proprietary and open-source models, and observe that restricted perceptual fields play a significant role in enabling active perception. Results reveal a significant gap in the active perception capability of MLLMs, indicating that this area deserves more attention. We hope that ActiView could help develop methods for MLLMs to understand multimodal inputs in more natural and holistic ways. [^1] Codes and data will be available at https://github.com/THUNLP-MT/ActiView
ActiView: Evaluating Active Perception Ability for Multimodal Large Language Models
Ziyue Wang*,1, Chi Chen*,1, Fuwen Luo*,1, Yurui Dong3, Yuanchi Zhang1, Yuzhuang Xu1, Xiaolong Wang1, Peng Li2, Yang Liu1,2 1Dept. of Comp. Sci. & Tech., Institute for AI, Tsinghua University, Beijing, China 2Institute for AI Industry Research (AIR), Tsinghua University, Beijing, China 3School of Management, Fudan University, Shanghai, China
@@ 1 Introduction
The advent of Multimodal Large Language Models (MLLMs) has marked a significant milestone in the realm of artificial intelligence, demonstrating capabilities that are increasingly approaching human-like performance (OpenAI, 2023; Liu et al., 2023c; Ye et al., 2024b). This advancement, while promising, also presents new challenges and opportunities for evaluating these models. As a result, the landscape of MLLM evaluation is rapidly evolving, with numerous benchmarks being developed to either comprehensively evaluate models (Fu et al., 2023; Liu et al., 2023d) or to analyze specific aspects of their capabilities (Liu et al., 2023a; Lu et al., 2023; Luo et al., 2024; Xiao et al., 2024; Li et al., 2024b; Nie et al., 2024; Qian et al., 2024).
@@
Figure 1: Active perception allows humans or models to perform more complex tasks by actively seeking and processing relevant information. In this paper, we evaluate two key active perception abilities for MLLMs: 1) shifting, as real-world scenarios often present limited views and require shifts to obtain new perspectives, and 2) zooming, which helps enhance perception by zooming out for a broader view and zooming in for details.
@@
General MME (Fu et al., 2023) Visual comprehension 1.1k 1.3k Manual
@@
Specialized ViP-Bench (Biernacki et al., 2021) Understanding of visual prompt 303 303 Manual
@@
V* (Wu and Xie, 2023) Fine-grained visual search 191 191 Manual
@@
Table 1: Comparison with other benchmarks for MLLMs. “Per. Fields”: Perceptual Fields. 1.9k*: Videos. Manual*: A mixture of manual annotation and data from existing benchmarks. Our benchmark focuses on evaluating active perception abilities via changes in visual perceptual fields, including shifting for compensating missing information, zooming for fine-grained details in the current fields, and a combination of both to mimic real-world scenarios.
Despite the extensive efforts devoted to MLLM evaluation, active perception (Bajcsy, 1988; Bajcsy et al., 2018) remains underexplored. Active perception involves understanding the reasons for sensing, choosing what to perceive, and determining the methods, timing, and locations for achieving that perception (Bajcsy et al., 2018). This is important because in the real world, the desired information often does not appear directly in the center of one’s field of vision. Instead, it requires individuals to move their field of view, locate details, and filter out distracting information. For example, in Figure 1, suppose we are looking for information in a giant painting. We need to first shift our view to locate the specific area and then possibly zoom in to gather detailed information. Intuitively, active perception not only enables a person or model to accomplish more complex tasks, but also has the potential to serve as a good indicator of the level of intelligence of a model, making it a critical capability that warrants thorough evaluation.
However, existing multimodal evaluation benchmarks are not well-suited to assess active perception capabilities. Table 1 summarizes several widely used or recently proposed multimodal evaluation benchmarks, most of them assess models in static perceptual field settings, where models process information presented directly to them without requiring active exploration or dynamic adjustments to their field of view. BLINK (Fu et al., 2024b), V* (Wu and Xie, 2023), and CNT (Roberts et al., 2023) are exceptions, as they utilize dynamic perceptual fields. However, they only consider either shifting or zooming of the field of view in specific scenarios, which are insufficient for measuring active perception capabilities. Therefore, there is a clear need for new evaluation frameworks that can adequately capture active perception abilities across diverse and dynamic environments.
To fill this gap, we introduce a novel benchmark specifically designed to evaluate Active perception through View changes (ActiView). Given the difficulty of comprehensively evaluating such capabilities across all possible scenarios, ActiView focuses on a series of tasks that are feasible to evaluate, yet still present significant challenges to current models. We manually curate a diverse set of instances, each including question-answer pairs and reasoning clues, and follows the Visual Question Answering (VQA) (Antol et al., 2015) format but exhibits additional features: 1) Each question requires an understanding of multiple detailed visual clues in the image to answer accurately. 2) View constraints are imposed, allowing models to perceive only a partial field of view of the full image at a time. This setup explicitly requires models to perform view shifting and zooming to gather necessary information and eliminate potential distractions, simulating the active perception process in real life. 3) In addition to answering visual questions, intermediate reasoning behaviors, such as view selection, also contribute to the evaluation.
Results from over 30 models reveal that these models generally lag behind in active perception. For instance, the strong proprietary model, GPT-4o, only achieved an average score of 66.40% with our designed evaluation pipelines for fundamental abilities, which is notably lower then the human score of 84.67%. Regarding another pipeline that allows models to flexibly integrate these fundamental abilities, GPT-4o achieves 69.54%, implying that combining fundamental active perception abilities can contribute to improvements. Moreover, the average performance gap between proprietary models and open-source models in active perception is considerably smaller within our designed pipelines than those observed in tasks from previous research. Recent small open-source models, in particular, exhibit approaching GPT-4o results. Experimental results suggest that models tend to perform better when given a complete image but struggle to develop a holistic understanding when presented with even all the separate and constrained perceptual fields. These findings highlight the need for further research in active perception and the value of our benchmark for advancing this field.
Figure 2: Examples of ActiView, exhibiting the following features: i) requiring focusing on multiple fine-grained regions; ii) requiring distinguishing distracting information from the entire image; iii) requiring moving of perceptual fields to obtain sufficient visual information to answer questions. During evaluation, models will be given an initial view cropped from the original image as shown above. Visual Information: human-annotated visual clues.
2 Related Works
@@ 2.1 MLLM Benchmarks
Extensive efforts have been devoted to developing MLLM evaluation benchmarks (Table 1). covering a wide range of capabilities, including visual comprehension  (Fu et al., 2023; Liu et al., 2023d; Fu et al., 2024a), visual perception (Fu et al., 2024b), hallucination (Liu et al., 2023a), and mathematical and logical reasoning (Lu et al., 2023; Xiao et al., 2024). However, most of them rely on a static view of the input image, which is not suited for assessing active perception. While BLINK (Fu et al., 2024b) involves view shifting, and both V* (Wu and Xie, 2023) and CNT (Roberts et al., 2023) require view zooming, active perception is not a prerequisite for solving their evaluation questions, making them insufficient for comprehensive active perception evaluation. In contrast, our benchmark considers both view shifting and zooming, with questions specifically designed to necessitate active perception for answering, which makes it a more robust framework towards active perception evaluation.
@@ 2.2 Active Perception in MLLMs
Although MLLMs have attracted extensive interest, less effort has been dedicated to improving the active perception capability of MLLMs. One line of research focuses on improving the ability of processing high-resolution images by using higher-resolution ViTs (Ye et al., 2024b), slicing high-resolution images and then concatenate them (Liu et al., 2024), or directly using LLMs to process raw patches of any resolution (Li et al., 2023a). The other line emphasizes visual search for fine-grained details. SEAL (Wu and Xie, 2023) fine-tunes a framework of two MLLMs to follow the visual search mechanism for precise visual grounding, and V-IRL (Yang et al., 2024) proposes an active detection strategy to improve the comprehension of real-world geospatial information. Despite these efforts, our evaluation results reveal that existing MLLMs still generally lack active perception capabilities. Our benchmark will shed light on evaluating and enhancing active perception in MLLMs.
@@ 3 ActiView
Our benchmark examines active perception abilities through different perceptual fields, where Actively zooming and shifting of Views (ActiView) are required. We summarize zooming and shifting as core components of active perception, as depicted in Figure 1, which allow us to evaluate active perception abilities of models both separately and integratedly. ActiView imitates the behavior of active perception by providing models with a constraint initial view, either a cropped field of the original image or a full image at reduced resolution. As shown in Figure 2, models should search for missing critical information through view zooming and shifting, while eliminating distractions caused by redundant content within the view.
@@ 3.1 Benchmark Overview
When perceiving an image, humans instinctively focus on three principle aspects: the depicted environment, the primary objects, and the events in which these objects are involved. Similarly, we categorize questions in our benchmark into three main types, as shown in Figure 2, which are further divided into eight sub-classes according to the type of visual information and features required to answer the questions. Due to page limitations, detailed descriptions and typical examples for each sub-class are provided in Appendix A.3. Below is a concise overview of categories in ActiView:
  • Environment-centric (Type I) involves three sub-classes. Geo-Localization (Geo-Loc) focuses on geographical features unique to specific countries or cities, requiring models to identify geographical locations implied in the image. Orientation (Orient) challenges models to reason from information of natural orientation. Daily-location (Daily-Loc) distincts from Geo-Loc by centering on everyday locations that could appear in most cities, without being tied to a particular country or city.
  • Object-centric (Type II) tasks go beyond simple grounding that directly ask for the attributes or relations of objects, by challenging models with distracting information. Three sub-classes are Object-attribute (Obj-Attr), that focuses on identifying objects attributes out of distractions that potentially mislead the model; Object-relation (Obj-Rel), that examines spatial relationships among objects while the questions do not explicitly ask for them; and Counting (Count-Dis), that involves counting objects while handling similar but distracting elements that can lead to incorrect answers.
  • Event-centric (Type III) focuses on the interactions between humans and objects, such as actions and activities. This category is divided according to the number of objects involved in the target event, including Event-single (Event-S) which focuses on events involving a single item or person; and Event-multi (Event-M) that happens among multiple items or people.
Figure 3: The statistical distribution of our benchmark.
Figure 4: Evaluation pipelines as described in §E. (a) Zooming requires models to select multiple regions to zoom in. It tests one of the fundamental active perception abilities. (b) Shifting challenges models to ask for more necessary information. It tests the other fundamental active perception abilities. (c) Mixed simulates human behavior when shifting perceptual fields for missing information. It is more flexible and applicable in real life compare to the previous two fundamental abilities. Note that while we provide an example in the figure where model delete a zoomed sub-view, the deletion behavior is NOT required. It is to address the compound features of the mixed pipeline (c) compare to the other fundamental pipelines (a) and (b).
3.2 Data Curation and Statistics
Our dataset is manually curated, including image collection, question and option annotation, and visual clue identification. To assess active perception abilities, which require zooming for fine-grained details and shifting for capturing missing information, we select images featuring multiple fine-grained objects and complex scenes or events. Detailed descriptions of the collected images and annotation guidelines can be found in Appendix A. Annotators are also required to identify visual clues to support their answers, as shown in the “Visual Information” columns in Figure 2. To prevent models from selecting answers based solely on the provided options, we adopt a set of more flexible annotation rules than the typical two- or four-option format. Option count ranges from two to seven, many of which are derived directly from the images. For instance, in the Geo-Loc task shown in Figure 2, the options all correspond to visual clues in the image, such as flags representing the UK, France, Spain, and Hungary. Furthermore, for options comprised by numbers, they are arranged in random order to avoid biased predictions.
The experiments of automatic data generation are discussed in Appendix K. In summary, we found that powerful models, such as GPT-4V and GPT-4o, fail to satisfy our annotation guidelines. They struggle with hallucination when processing multiple images, and fall short on distinguish between visual facts in the image and external world knowledge not present in the image.
Statistics.
We collected 314 images and annotated 325 questions, with distribution of categories shown in Figure 3. To further enrich the diversity, each question corresponds to 5 different evaluation instances, assessing active perception across varying components and difficulties. In total, 1,625 evaluation instances are curated. On average, there are 3.24 options and 2.64 sub-views containing visual clues per question, highlighting that a single view is often insufficient for answering accurately, and that the ability to comprehend multiple images jointly is crucial for our benchmark.